OpenStudy (anonymous):

Help! If you deposit $750 in an account that pays 8% annual interest compounded continuously, what will the balance be after five years? Explain please?

OpenStudy (anonymous):

Use A= P(1+r/n)^nt

OpenStudy (anonymous):

Compound Interest Formula

OpenStudy (anonymous):

what do the variables stand for?

OpenStudy (anonymous):

P=initial dollar amount; r=interest rate expressed as a decimal fraction t=time (in years)

OpenStudy (anonymous):

So.. \[A = 750(1+\frac{ 0.08 }{ 8 })^{8(5)}\]

OpenStudy (anonymous):

Give me the answer choices

OpenStudy (anonymous):

oh okay! also the answer choices are a. 1,277.16 b. 1,118.87 c. 1,832.06 d. 29,803.91

OpenStudy (anonymous):

oops the n is number of years.. Let me do it again

OpenStudy (anonymous):

So we have a continuous compounding... For continuous compounding, A=Pe^(rt), where P=initial dollar amount; r=interest rate expressed as a decimal fraction; and t=time (in years). Is this sufficient info to enable you to calculate A (the amount the account owner will have in the account after 5 years?

OpenStudy (anonymous):

yes! P=750 R=8% T=5